OpenStudy (jordanloveangel):

A line has a slope of 3 and contains the point (−1, −8). Which equations represent the line? Choose all answers that are correct. A. −3x + y = −5 B. 3x + y = −5 C. (y – 8) = 3(x − 1) D. (y + 8) = 3(x + 1)

OpenStudy (dtan5457):

Put this into point slope form first. (y-y1)=m(x-x1)

OpenStudy (anonymous):

Just about to say that :)

OpenStudy (dtan5457):

D. (y + 8) = 3(x + 1) will definitely one of them, now distribute the parenthesis and see what other answers work.

OpenStudy (dtan5457):

IF they work at all..

OpenStudy (mrnood):

a straight line always has an equation in the form y=mx+b where m is the slope so in your case y = 3x +b and you know that when x=-1 then y= -8 so you can find b from that equation

OpenStudy (jordanloveangel):

im thinking the answer is d but i m never sure

OpenStudy (anonymous):

The definite answer is D.

OpenStudy (dtan5457):

There is one more I believe

OpenStudy (jordanloveangel):

okay thx guys

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

Hold on.......

OpenStudy (anonymous):

Stay put..

OpenStudy (jordanloveangel):

it says Choose all answers that are correct.

OpenStudy (dtan5457):

If you distribute out (y + 8) = 3(x + 1) you get y+8=3x+3

OpenStudy (dtan5457):

One other answer represents that.

OpenStudy (dtan5457):

Not in the exact form

OpenStudy (mrnood):

the correct ANSWER is use the above formula and work it out yourself. IF you take other people#'s direct answers you cannot know they are right, and you learn nothing

OpenStudy (dtan5457):

I'll give you a hint. y+8=3x+3 Isolate the x and y's to one side, and the whole numbers to another

OpenStudy (anonymous):

Ha. yep.

OpenStudy (jordanloveangel):

wait im thinking the right answesr D C and A right?

OpenStudy (anonymous):

exactly, you have to rearrange the equation. what will you get?

OpenStudy (dtan5457):

It can't be C. C. (y – 8) = 3(x − 1) That's just the incorrect way to plug into the point slope formula.

OpenStudy (anonymous):

Don't guess. Math is not a guessing subject. There are always definite answers.

OpenStudy (jordanloveangel):

so AB and D?

OpenStudy (mrnood):

@jordanloveangel it CANNOT be C & D - they have different sign for +/- 8 in th esdame form of equaiton

OpenStudy (jordanloveangel):

this is so confusing

OpenStudy (mrnood):

Not if YOU work through it methodically, instead of reading all th eposts and guessing

OpenStudy (dtan5457):

@MrNood It can't be D?

OpenStudy (jordanloveangel):

im so condusing oh how 2 slove and get the answer

OpenStudy (mrnood):

I said it cn't be C AND D, not thast it can't be C OR D

OpenStudy (dtan5457):

How is not D?

OpenStudy (jordanloveangel):

??????

OpenStudy (mrnood):

I didn't say that it's not D the OP guessed a,c AND D I said that it cannot be BOTH C and D because they are conflicting

OpenStudy (dtan5457):

Oh, that clears it up.

OpenStudy (jordanloveangel):

okay then wuts the right answers

OpenStudy (jordanloveangel):

@paki

OpenStudy (mrnood):

as I said above - the right answer is 'use the formula to work it out' anything else is someone else's answer and you cannot know if they have just guessed like oyu

OpenStudy (jordanloveangel):

@AlexandervonHumboldt2 can u help me plase

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

y=-1/2x+3 like i said earlier B C D

OpenStudy (alexandervonhumboldt2):

you posted the same queston

OpenStudy (jordanloveangel):

wait wrong question

OpenStudy (jordanloveangel):

A line has a slope of 3 and contains the point (−1, −8). Which equations represent the line? Choose all answers that are correct. A. −3x + y = −5 B. 3x + y = −5 C. (y – 8) = 3(x − 1) D. (y + 8) = 3(x + 1)

OpenStudy (jordanloveangel):

this is my question

OpenStudy (alexandervonhumboldt2):

so -8=3*-1+b b=11 y=3x+11 slope intersept let me finish

OpenStudy (alexandervonhumboldt2):

so convert ech other equation and chck if othrs will b eual to this

OpenStudy (jordanloveangel):

okay so is the answer A ,BandD

OpenStudy (alexandervonhumboldt2):

yes should be

OpenStudy (jordanloveangel):

kk thank yoou verrrryyyyyyyyyyyyy much

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (mrnood):

@jordanloveangel that is NOT coorect

OpenStudy (jordanloveangel):

dude when wut the hell is the right dam answer

OpenStudy (mrnood):

look at B 3x+y=-5 the point x=-1, y=-8 is on the line if oyu put x=-1 and y=-8 into equation B it is not correct

OpenStudy (anonymous):

Wouldn't it be -3+y=-5, because one you have the equation after distributing, you move the 3x.

OpenStudy (mrnood):

@csheveland WHAT?

OpenStudy (anonymous):

sorry -3x+y=-5

OpenStudy (mrnood):

@jordanloveangel Surely you can see from th efiasco above that the BEST ooption is to understand the problem and solve it for yourself. Do you want help with equations of straight lines?

OpenStudy (jordanloveangel):

nvm i got the answers

OpenStudy (mrnood):

@jordanloveangel and what do you NOW think are the correct ones (there are 2 only...)

OpenStudy (jordanloveangel):

the right answers r a and d

OpenStudy (mrnood):

@AlexandervonHumboldt2

OpenStudy (jordanloveangel):

@Loser66

OpenStudy (loser66):

A

OpenStudy (jordanloveangel):

@Loser66 there r 2 right answers

OpenStudy (loser66):

D

OpenStudy (jordanloveangel):

thx